# Disc3D: Automatic Curation of High-Quality 3D Dialog Data via Discriminative Object Referring

### v2

@@ Abstract
3D Multi-modal Large Language Models (MLLMs) still lag behind their 2D peers, largely because large-scale, high-quality 3D scene–dialogue datasets remain scarce. Prior efforts hinge on expensive human annotation and leave two key ambiguities unresolved: viewpoint ambiguity—spatial language presumes unknown camera poses, and object referring ambiguity—non-exclusive descriptions blur the line between targets and distractors. We therefore present a fully automated pipeline that converts raw 3D scans into unambiguous, high-quality dialogue data at a fraction of the previous cost. By synergizing rule-based constraints with 2D MLLMs and LLMs, the pipeline enables controllable, scalable generation without human intervention. The pipeline comprises four stages: (1) meta-annotation collection harvesting object-, frame-, and scene-level captions, (2) scene graph construction with relation correction to capture the proximal object relations, (3) discriminative object referring that generates exclusive and compact descriptions, and (4) multi-task data generation synthesizing diverse dialogues. Our pipeline systematically mitigates inherent flaws in source datasets and produces the final Disc3D dataset—over 2 million samples in 25K hybrid 3D scenes, spanning scene&view&object captioning, visual grounding, and five object-centric QA tasks. Extensive experiments demonstrate that training with Disc3D yields consistent, significant improvements on both public benchmarks and our multifaceted Disc3D-QA tasks. Code, data, and models will be publicly available.
@@ 1 Introduction
Existing 3D MLLM datasets predominantly rely on manual annotation [lyu2024mmscan, azuma2022scanqa, ma2022sqa3d, achlioptas2020referit3d, chen2020scanrefer, zhang2023multi3drefer, yang2025thinking], which is inherently costly and limits scalability. Moreover, free-form annotations introduce two critical ambiguities: 1)Viewpoint ambiguity: As Fig. 1a shows, spatial descriptions often presuppose a specific 2D camera pose that is unknown during training and evaluation; 2) Object referring ambiguity: The unconstrained object referring expressions lack discriminative features, making target objects indistinguishable from distractors, as shown in Fig. 1b. Meanwhile, as Fig. 2 shows, inherent deficiencies in existing 3D scene datasets substantially increase the difficulty of multi-modal data generation. These issues necessitate an automated, systematic, scalable pipeline to convert raw scans into unambiguous, high-quality 3D dialogue data.
@@ 2 Related Work
Early work exploited 3D scene scans [dai2017scannet, wald2019rio, chang2017matterport3d, procthor, baruch2021arkitscenes, ramakrishnan2021hm3d, mao2022multiscan] to establish diverse 3D multi-modal benchmarks, including 3D question answering (QA) [azuma2022scanqa, ma2022sqa3d, yang2025thinking, OpenEQA2023] and 3D visual grounding [chen2020scanrefer, achlioptas2020referit3d, zhang2023multi3drefer]. 3D QA task assesses 3D MLLM’s understanding of objects and layouts in the scene [zhu20233d, chen2024ll3da, huang2024chat], while 3D visual grounding tasks evaluate its ability to localize objects  [yuan2021instancerefer, yang2021sat, huang2021text, luo20223d, wu2023eda, xu2024vlm, chen2023unit3d]. Motivated by advances in Large Language Models (LLMs) and 2D Multi-modal LLMs (MLLMs), more recent efforts [jia2024sceneverse, lyu2024mmscan] combine human annotations with 2D MLLM capabilities to generate large-scale 3D scene data. However, these pipelines suffer from viewpoint ambiguity and referential uncertainty due to unconstrained annotations, and incur high production costs because of extensive manual labeling. We address these challenges with a fully automated pipeline based on Discriminative Object Referring, producing a larger, more diverse corpus free from previous ambiguities and cost constraints.
@@
Building on the rapid progress of 2D MLLMs [bai2025qwen2, zhu2024llava, wu2024deepseek, team2025Kimi], recent research integrates 3D visual information into LLMs to create 3D MLLMs. Proposed methods include point‑cloud modules [avetisyan2024scenescript, deng20253d, xu2024pointllm, qi2024gpt4point], lifted 2D foundation features [zhu2024llava], video‑based representations [zheng2025video], instance‑level embeddings [yu2025inst3d], and bird’s‑eye‑view projections [GPT4Scene]. Owing to the increased dimensionality and complexity of 3D data, no single architecture yet demonstrates robust spatial understanding from limited training samples.
@@ 3.2 Deficiencies in Source Data
Because the label vocabularies differ across source corpora, we let a strong LLM [liu2024deepseek] build a single directed graph that encodes the subsumption structure. Each node is an object label; a directed edge $u\rightarrow v$ exists whenever LLM confirms that $v$ is a subtype of $u$ (Fig. 2a). Candidate edges are first shortlisted by word overlap and then validated by LLM in a zero-shot fashion.
@@ 3.3 Data Curation Pipeline
For scene acquisition, we compile 25 K scans from five datasets: ScanNet [dai2017scannet], CA-1M [lazarow2025cubify], 3RScan [wald2019rio], Structured3D [zheng2020structured3d] (training), and ScanNet++ [yeshwanthliu2023scannetpp] (testing), covering both real and synthetic environments. The subsequent curation pipeline consists of four stages.
@@
  • Caption annotation. For scenes & frames, we prompt MLLM with given images and instructions, while max-coverage sampling [zheng2025video] caps the number of frames used for the scene caption. For objects, we rank viewpoints by (i) the 2D distance from the object center to the image center and (ii) the visible area ratio obtained by rendering the annotated mesh. The top-2 views are cropped and sent to the MLLM for captioning; an LLM then extracts appearance attributes.
@@
Following SceneVerse [jia2024sceneverse], we approximate object occupancy with 3D bounding boxes and build an object-centric graph that records only vertical adjacency and horizontal contact. To suppress noise and ambiguity, we discard multi-object relations (involving more than two nodes) and all relative directions (e.g., left/right or clock-face positions).
Rule-based heuristics cannot identify subtle cases such as “inside” or “hanging from”. We therefore introduce a Relation Correction stage, illustrated in Fig. 4a. An MLLM first narrates the interaction between co-visible objects in free text; an R1-style LLM [guo2025deepseek] then parses each sentence into one of the predefined relation templates, updating the initial graph accordingly.
@@
Figure 5: Discriminative object referral examples in Scannet++ [yeshwanthliu2023scannetpp] scans.
@@
SceneVerse [jia2024sceneverse] 2.5M 68K No No No
MMScan [lyu2024mmscan] 1.45M 5.2K No No No ★ ★ ✩
@@
Table 1: Comparison with two other large-scale 3D MLLM multi-task datasets. Disc3D autonomously curates a two-million-sample multi-task corpus by means of Discriminative Object Referencing, a carefully engineered module that injects contextual constraints into object descriptions and eliminates ambiguities in object identification and spatial relations. Earlier 3D vision-language datasets [azuma2022scanqa, ma2022sqa3d, achlioptas2020referit3d, zhang2023multi3drefer], being smaller and task-specific, still remain vulnerable to ambiguity; they are therefore excluded from direct comparison.
@@ 3.4 Statistics & Analysis
MMScan [lyu2024mmscan] and SceneVerse [jia2024sceneverse] assemble large-scale 3D multi-modal datasets with human-in-the-loop pipelines. Disc3D eliminates manual effort by replacing annotators with a rule-based engine powered by MLLMs and LLMs. The resulting corpus spans more scenes and samples than MMScan and subsumes the task diversity of SceneVerse. Like MMScan, Disc3D provides scene- and region-level captioning, visual grounding, and multi-dimensional QA. We extend this set with view-level captioning derived from individual frames. Throughout data curation, we eliminated viewpoint ambiguities introduced by either manual design or LLM-generated text. To resolve object-referring ambiguity, we propose Comparative Disambiguation and Spatial Anchoring, which embed discriminative constraints into object descriptions. A brief comparison is provided in Tab. 1.
@@
In total, 12% of the descriptions ($130/1092$) were revised. As illustrated in Fig. 7, the majority of edits introduced by human annotators do not rectify factual inaccuracies but instead refine imprecise wording. For instance, limited descriptors such as “largest” or “smallest” often inadequately capture the distinctive size characteristics of irregularly shaped objects (e.g., floor lamps) or soft, deformable items (e.g., blankets). We contend that such imprecisions, though potentially influencing model preferences, are comparatively tolerable relative to factual errors or ambiguous expressions. Given that manually annotated benchmarks still exhibit non-negligible noise—e.g., human re-evaluation in [huang2025unveiling] reports only 62% and 80% accuracy on ScanQA and SQA3D, respectively—Disc3D provides 3D multi-modal data with comparatively low label noise.
@@
3D MLLM Ross3D [wang2025ross3d] 0.106 0.672 0.529 0.181 0.232 0.308 / – 0.630 / 0.657
Video-3DLLM [zheng2025video] 0.150 0.638 0.211 0.002 0.188 0.301 / – 0.586 / –
LLaVA-3D [zhu2024llava] 0.071 0.590 0.514 0.046 0.580 – / 0.306 – / 0.601
@@ 4 Experiments
We evaluate on two canonical 3D scene-understanding tasks. (1) 3D visual grounding localises the object described by a natural-language utterance. Asking 3D MLLMs to regress a 3D bounding box in text space is unreliable. We therefore benchmark proposal-matching methods [zheng2025video, wang2025ross3d] that recast the task as proposal classification: the model aligns predicted grounding tokens with 3D proposals extracted by an off-the-shelf detector [schult2022mask3d] at the feature level. We report accuracy on the Disc3D grounding split with 2790 samples and analyse performance under varying conditions. (2) 3D question answering uses a linear, objective metric that cleanly reflects different facets of model capability; consequently, our main experiments and ablations centre on QA. Benchmarks include the public datasets [azuma2022scanqa, ma2022sqa3d] and the Disc3D multi-dimensional QA suite. The Disc3D-QA test set contains 1 000 questions per category across five dimensions; object counts are unavailable at training time, making this sub-task zero-shot for our model.
@@
For visual grounding, we report exact-match accuracy: we use ground-truth object boxes as proposals, thereby removing any confounding effect of localization error. For question-answering, we adopt top-1 exact-match (EM) and its relaxed variant (EM-R) for normal questions, and Mean Relative Accuracy [yang2025thinking] (MRA) for numerical questions such as object size and absolute distance.
@@
We re-implement LLaVA-3D [zhu2024llava] on InternVL3 [zhu2025internvl3] and fine-tune the MLP projector and LLM part. For our main experiments and ablations, leveraging its compact end-to-end design and competitive efficiency. In addition to Disc3D, we combine 220 K samples from public datasets to enlarge diversity and enable direct comparison with prior work (see Tab. 4). To balance cost and coverage we sample at most 32 frames per scene with max-coverage sampling [zheng2025video]; all images are centre-cropped to $336\times 336$ pixels.
@@
Video-3D-LLM [zheng2025video] Ross3D [wang2025ross3d]
@@ 6.2 Label Graph Construction
To construct the label graph, we adopt DeepSeek-V3 [liu2024deepseek] as the oracle for deciding subsumption between object labels. The query template is:
@@ 6.3 Details of Source Dataset Processing
    Class-agnostic annotations in CA-1M. CA-1M [lazarow2025cubify] provides object annotation without semantic labels. During the Object-Caption stage, we prompt a 2D MLLM with the object views and ask it to predict the class label, following the NYU-40 taxonomy.
@@
    Invalid scenes in Structured3D. We filter out the scenes listed in the official error list. We further discard any image whose mean intensity exceeds a brightness threshold, thereby removing over-exposed views. Specifically, for each image, we compute the fraction of uint8 pixels whose gray-level intensity exceeds 245. If this fraction surpasses 90%, the image is labeled as over-exposed. Following this criterion, we discard 143 over-exposed scenes from Structured3D [zheng2020structured3d]; representative examples are shown in Figure 9.
@@
    Missing segmentation labels in CA-1M. Because CA-1M [lazarow2025cubify] lacks 2D and 3D segmentation, we estimate object visibility masks used in the caption annotation procedure heuristically. Inspired by OpenYolo3D [boudjoghra2024open], we combine depth values with the provided 2D bounding boxes to construct a per-frame label map. For each image, we compute the average depth inside every bounding box, sort the boxes from far to near, and assign object IDs in this order. Consequently, regions of distant objects are covered by nearer ones.
@@ 6.10 Comparison of Object Descriptions
As shown in Figure 15, we provide a comparison of object descriptions. MMScan [lyu2024mmscan] and SceneVerse [jia2024sceneverse] enumerate almost every observable attribute — appearance, location, geometry, functionality, and ancillary details, yielding exhaustive yet lengthy descriptions. Such verbosity, while informative, dilutes the discriminative cues required to unambiguously identify an object. In contrast, our Disc3D pipeline distills a concise, multi-view signature that foregrounds the distinguishing characteristics of each object, thereby offering a more effective dataset for assessing and guiding the spatial understanding capability of 3D MLLMs.

### v3

@@ Abstract
3D Multi-modal Large Language Models (MLLMs) still lag behind their 2D peers, largely because large-scale, high-quality 3D scene–dialogue datasets remain scarce. Prior efforts hinge on expensive human annotation and leave two key ambiguities unresolved: viewpoint ambiguity—spatial language presumes unknown camera poses, and object referring ambiguity—non-exclusive descriptions blur the line between targets and distractors. We therefore present a fully automated pipeline that converts raw 3D scans into unambiguous, high-quality dialogue data at a fraction of the previous cost. By synergizing rule-based constraints with 2D MLLMs and LLMs, the pipeline enables controllable, scalable generation without human intervention. The pipeline comprises four stages: (1) meta-annotation collection harvesting object-, frame-, and scene-level captions, (2) scene graph construction with relation correction to capture the proximal object relations, (3) discriminative object referring that generates exclusive and compact descriptions, and (4) multi-task data generation synthesizing diverse dialogues. Our pipeline systematically mitigates inherent flaws in source datasets and produces the final Disc3D dataset—over 2 million samples in 25K hybrid 3D scenes, spanning scene&view&object captioning, visual grounding, and five object-centric QA tasks. Extensive experiments demonstrate that training with Disc3D yields consistent, significant improvements on both public benchmarks and our multifaceted Disc3D-QA tasks. Our codebase is available at https://github.com/bytedance/Disc3D, and the dataset at https://huggingface.co/datasets/Sywwwwww/Disc3D.
@@ 1 Introduction
Existing 3D MLLM datasets predominantly rely on manual annotation [22, 3, 23, 1, 8, 42, 37], which is inherently costly and limits scalability. Moreover, free-form annotations introduce two critical ambiguities: 1)Viewpoint ambiguity: As Fig. 1a shows, spatial descriptions often presuppose a specific 2D camera pose that is unknown during training and evaluation; 2) Object referring ambiguity: The unconstrained object referring expressions lack discriminative features, making target objects indistinguishable from distractors, as shown in Fig. 1b. Meanwhile, as Fig. 2 shows, inherent deficiencies in existing 3D scene datasets substantially increase the difficulty of multi-modal data generation. These issues necessitate an automated, systematic, scalable pipeline to convert raw scans into unambiguous, high-quality 3D dialogue data.
@@ 2 Related Work
Early work exploited 3D scene scans [11, 31, 7, 12, 5, 28, 25] to establish diverse 3D multi-modal benchmarks, including 3D question answering (QA) [3, 23, 37, 24] and 3D visual grounding [8, 1, 42]. 3D QA task assesses 3D MLLM’s understanding of objects and layouts in the scene [47, 9, 15], while 3D visual grounding tasks evaluate its ability to localize objects  [41, 38, 17, 21, 33, 35, 10]. Motivated by advances in Large Language Models (LLMs) and 2D Multi-modal LLMs (MLLMs), more recent efforts [18, 22] combine human annotations with 2D MLLM capabilities to generate large-scale 3D scene data. However, these pipelines suffer from viewpoint ambiguity and referential uncertainty due to unconstrained annotations, and incur high production costs because of extensive manual labeling. We address these challenges with a fully automated pipeline based on Discriminative Object Referring, producing a larger, more diverse corpus free from previous ambiguities and cost constraints.
@@
Building on the rapid progress of 2D MLLMs [4, 45, 34, 30], recent research integrates 3D visual information into LLMs to create 3D MLLMs. Proposed methods include point‑cloud modules [2, 13, 36, 26], lifted 2D foundation features [45], video‑based representations [43], instance‑level embeddings [40], and bird’s‑eye‑view projections [27]. Owing to the increased dimensionality and complexity of 3D data, no single architecture yet demonstrates robust spatial understanding from limited training samples.
@@ 3.2 Deficiencies in Source Data
Because the label vocabularies differ across source corpora, we let a strong LLM [20] build a single directed graph that encodes the subsumption structure. Each node is an object label; a directed edge $u\rightarrow v$ exists whenever LLM confirms that $v$ is a subtype of $u$ (Fig. 2a). Candidate edges are first shortlisted by word overlap and then validated by LLM in a zero-shot fashion.
@@ 3.3 Data Curation Pipeline
For scene acquisition, we compile 25 K scans from five datasets: ScanNet [11], CA-1M [19], 3RScan [31], Structured3D [44] (training), and ScanNet++ [39] (testing), covering both real and synthetic environments. The subsequent curation pipeline consists of four stages.
@@
  • Caption annotation. For scenes & frames, we prompt MLLM with given images and instructions, while max-coverage sampling [43] caps the number of frames used for the scene caption. For objects, we rank viewpoints by (i) the 2D distance from the object center to the image center and (ii) the visible area ratio obtained by rendering the annotated mesh. The top-2 views are cropped and sent to the MLLM for captioning; an LLM then extracts appearance attributes.
@@
Following SceneVerse [18], we approximate object occupancy with 3D bounding boxes and build an object-centric graph that records only vertical adjacency and horizontal contact. To suppress noise and ambiguity, we discard multi-object relations (involving more than two nodes) and all relative directions (e.g., left/right or clock-face positions).
Rule-based heuristics cannot identify subtle cases such as “inside” or “hanging from”. We therefore introduce a Relation Correction stage, illustrated in Fig. 4a. An MLLM first narrates the interaction between co-visible objects in free text; an R1-style LLM [14] then parses each sentence into one of the predefined relation templates, updating the initial graph accordingly.
@@
Figure 5: Discriminative object referral examples in Scannet++ [39] scans.
@@
SceneVerse [18] 2.5M 68K No No No
MMScan [22] 1.45M 5.2K No No No ★ ★ ✩
@@
Table 1: Comparison with two other large-scale 3D MLLM multi-task datasets. Disc3D autonomously curates a two-million-sample multi-task corpus by means of Discriminative Object Referencing, a carefully engineered module that injects contextual constraints into object descriptions and eliminates ambiguities in object identification and spatial relations. Earlier 3D vision-language datasets [3, 23, 1, 42], being smaller and task-specific, still remain vulnerable to ambiguity; they are therefore excluded from direct comparison.
@@ 3.4 Statistics & Analysis
MMScan [22] and SceneVerse [18] assemble large-scale 3D multi-modal datasets with human-in-the-loop pipelines. Disc3D eliminates manual effort by replacing annotators with a rule-based engine powered by MLLMs and LLMs. The resulting corpus spans more scenes and samples than MMScan and subsumes the task diversity of SceneVerse. Like MMScan, Disc3D provides scene- and region-level captioning, visual grounding, and multi-dimensional QA. We extend this set with view-level captioning derived from individual frames. Throughout data curation, we eliminated viewpoint ambiguities introduced by either manual design or LLM-generated text. To resolve object-referring ambiguity, we propose Comparative Disambiguation and Spatial Anchoring, which embed discriminative constraints into object descriptions. A brief comparison is provided in Tab. 1.
@@
In total, 12% of the descriptions ($130/1092$) were revised. As illustrated in Fig. 7, the majority of edits introduced by human annotators do not rectify factual inaccuracies but instead refine imprecise wording. For instance, limited descriptors such as “largest” or “smallest” often inadequately capture the distinctive size characteristics of irregularly shaped objects (e.g., floor lamps) or soft, deformable items (e.g., blankets). We contend that such imprecisions, though potentially influencing model preferences, are comparatively tolerable relative to factual errors or ambiguous expressions. Given that manually annotated benchmarks still exhibit non-negligible noise—e.g., human re-evaluation in [16] reports only 62% and 80% accuracy on ScanQA and SQA3D, respectively—Disc3D provides 3D multi-modal data with comparatively low label noise.
@@
3D MLLM Ross3D [32] 0.106 0.672 0.529 0.181 0.232 0.308 / – 0.630 / 0.657
Video-3DLLM [43] 0.150 0.638 0.211 0.002 0.188 0.301 / – 0.586 / –
LLaVA-3D [45] 0.071 0.590 0.514 0.046 0.580 – / 0.306 – / 0.601
@@ 4 Experiments
We evaluate on two canonical 3D scene-understanding tasks. (1) 3D visual grounding localises the object described by a natural-language utterance. Asking 3D MLLMs to regress a 3D bounding box in text space is unreliable. We therefore benchmark proposal-matching methods [43, 32] that recast the task as proposal classification: the model aligns predicted grounding tokens with 3D proposals extracted by an off-the-shelf detector [29] at the feature level. We report accuracy on the Disc3D grounding split with 2790 samples and analyse performance under varying conditions. (2) 3D question answering uses a linear, objective metric that cleanly reflects different facets of model capability; consequently, our main experiments and ablations centre on QA. Benchmarks include the public datasets [3, 23] and the Disc3D multi-dimensional QA suite. The Disc3D-QA test set contains 1 000 questions per category across five dimensions; object counts are unavailable at training time, making this sub-task zero-shot for our model.
@@
For visual grounding, we report exact-match accuracy: we use ground-truth object boxes as proposals, thereby removing any confounding effect of localization error. For question-answering, we adopt top-1 exact-match (EM) and its relaxed variant (EM-R) for normal questions, and Mean Relative Accuracy [37] (MRA) for numerical questions such as object size and absolute distance.
@@
We re-implement LLaVA-3D [45] on InternVL3 [46] and fine-tune the MLP projector and LLM part. For our main experiments and ablations, leveraging its compact end-to-end design and competitive efficiency. In addition to Disc3D, we combine 220 K samples from public datasets to enlarge diversity and enable direct comparison with prior work (see Tab. 4). To balance cost and coverage we sample at most 32 frames per scene with max-coverage sampling [43]; all images are centre-cropped to $336\times 336$ pixels.
@@
Video-3D-LLM [43] Ross3D [32]
@@ 6.2 Label Graph Construction
To construct the label graph, we adopt DeepSeek-V3 [20] as the oracle for deciding subsumption between object labels. The query template is:
@@ 6.3 Details of Source Dataset Processing
    Class-agnostic annotations in CA-1M. CA-1M [19] provides object annotation without semantic labels. During the Object-Caption stage, we prompt a 2D MLLM with the object views and ask it to predict the class label, following the NYU-40 taxonomy.
@@
    Invalid scenes in Structured3D. We filter out the scenes listed in the official error list. We further discard any image whose mean intensity exceeds a brightness threshold, thereby removing over-exposed views. Specifically, for each image, we compute the fraction of uint8 pixels whose gray-level intensity exceeds 245. If this fraction surpasses 90%, the image is labeled as over-exposed. Following this criterion, we discard 143 over-exposed scenes from Structured3D [44]; representative examples are shown in Figure 9.
@@
    Missing segmentation labels in CA-1M. Because CA-1M [19] lacks 2D and 3D segmentation, we estimate object visibility masks used in the caption annotation procedure heuristically. Inspired by OpenYolo3D [6], we combine depth values with the provided 2D bounding boxes to construct a per-frame label map. For each image, we compute the average depth inside every bounding box, sort the boxes from far to near, and assign object IDs in this order. Consequently, regions of distant objects are covered by nearer ones.
@@ 6.10 Comparison of Object Descriptions
As shown in Figure 15, we provide a comparison of object descriptions. MMScan [22] and SceneVerse [18] enumerate almost every observable attribute — appearance, location, geometry, functionality, and ancillary details, yielding exhaustive yet lengthy descriptions. Such verbosity, while informative, dilutes the discriminative cues required to unambiguously identify an object. In contrast, our Disc3D pipeline distills a concise, multi-view signature that foregrounds the distinguishing characteristics of each object, thereby offering a more effective dataset for assessing and guiding the spatial understanding capability of 3D MLLMs.